OpenStudy (ammarah):

Completing the square. Find the value of n such that each expression is a perfect square trinomial. x^-6x+n

OpenStudy (ammarah):

@ganeshie8

OpenStudy (ammarah):

@whpalmer4

OpenStudy (whpalmer4):

To complete the square, the constant term needs to be the square of one half of the coefficient of the \(x\) term.

OpenStudy (ammarah):

i dont understant please help me what is the cnstant term?

OpenStudy (whpalmer4):

The constant term is the one that doesn't have the variable... \[x^2+3x+5\]the constant term here is 5

OpenStudy (ammarah):

howd u get 3>?

OpenStudy (whpalmer4):

That's an example...

OpenStudy (ammarah):

ok so in the question 6 isnt a square

OpenStudy (whpalmer4):

You didn't read what I wrote quite correctly. Take the coefficient of the \(x\) term. Divide in half. Square that number. That is the number you need as the constant term to form a perfect square.

OpenStudy (whpalmer4):

Let's work backwards: a perfect square can be written as \[(x+a)^2 = (x+a)(x+a) = x^2 +2ax + a^2\] Now, you have \[x^2-6x+n\]\[x^2+2ax+a^2\] for those to be equal, \[-6x=2ax\]and\[n=a^2\]

OpenStudy (whpalmer4):

If we solve the first for \(a\): \[-6x=2ax\]\[-6=2a\]\[-3=a\] That's half of the coefficient of the \(x\) term, right? Now we square it: \((-3)^2 = 9\) \[(x-3)^2 = x^2 -6x + 9\]

OpenStudy (whpalmer4):

\(n=9\) makes \(x^2-6x+n\) a perfect square.

OpenStudy (ammarah):

ohh ok i get it now...

OpenStudy (whpalmer4):

Now, if you are completing the square in the context of an equation, you need to be careful that you don't break the equality! For example: \[x^2 + 4x = 0\]Hopefully, you can easily see that we need to add \(4\) to the left side of the equation to complete the square and rewrite it as \((x+2)^2\). However, we can't just add \(4\) to one side of the equation unless we either add \(4\) to the other side of the equation, or we subtract \(4\) somewhere else on the same side where we added it. The two are of course equivalent, and which you do is just a matter of convenience. \[(x^2+4x + 4) - 4 = 0\]\[(x+2)^2-4 = 0\] there I added and subtracted on the same side, and \(+4 - 4 = 0\) so everything is fine. I could also have done this: \[x^2 + 4x + 4 = 0 + 4\]\[(x+2)^2 = 4\] I think you should be able to see that is equivalent to the other, right?

OpenStudy (ammarah):

hmm yeah how about if its like thsi 2x^2-8x+n

OpenStudy (whpalmer4):

Well, there you need to factor out that leading 2. \[2x^2-8x+n\] is the same as \(2(x^2-4x+n/2)\) right? Complete the square on the stuff inside the parentheses. Or, if you like you can do it like this: \[ax^2+bx+c\]take half of \((b/a)\), square that, and that's your new constant term.

OpenStudy (ammarah):

so it would be +16?

OpenStudy (ammarah):

@whpalmer4

OpenStudy (whpalmer4):

Sorry, I didn't explain that quite correctly: you also have to divide that squared value by the value of \(a\).... \[2(x-2)^2 = 2(x^2-4x+4) = 2x^2-8x+8\]

OpenStudy (ammarah):

and then?

OpenStudy (whpalmer4):

So, the correct value was +8, not +16...

OpenStudy (ammarah):

ok thanks can u check out teh other problem i posted? please